### 12 Mar 2026 v1

# Geometric inequalities and the Alexandrov-Bakelman-Pucci technique

Abstract.
In this expository paper, we discuss a unified framework for proving various geometric inequalities, based on the so-called Alexandrov-Bakelman-Pucci technique. Examples include Cabré’s proof of the classical isoperimetric inequality in Euclidean space; the Fenchel-Willmore-Chen inequality for the mean curvature of a submanifold; the sharp version of the Michael-Simon Sobolev inequality for submanifolds; the sharp version of Ecker’s logarithmic Sobolev inequality for submanifolds; and the Sobolev inequality for complete manifolds with nonnegative Ricci curvature and Euclidean volume growth. Finally, we discuss a connection to the work of Heintze and Karcher on the volume of a tubular neighborhood of a hypersurface in a manifold with nonnegative Ricci curvature.
The author was supported by the National Science Foundation under grant DMS-2403981 and by the Simons Foundation.
1. The classical isoperimetric inequality and the Sobolev inequality
The isoperimetric inequality is one of the most important inequalities in differential geometry.
Theorem 1.1 (Isoperimetric inequality).
Let $D$ be a compact domain in $\mathbb{R}^{n}$ with smooth boundary. Then $|\partial D|\geq n\,|B^{n}|^{\frac{1}{n}}\,|D|^{\frac{n-1}{n}}$, where $B^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$.
The isoperimetric inequality is closely related to the Brunn-Minkowski inequality (see [25]) and to the Sobolev inequality.
Theorem 1.2 (Sobolev inequality).
Let $D$ be a compact domain in $\mathbb{R}^{n}$ with smooth boundary. Let $f$ be a smooth positive function on $D$. Then
$\int_{D}|\nabla f|+\int_{\partial D}f\geq n\,|B^{n}|^{\frac{1}{n}}\,\Big(\int_{D}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}},$
where $B^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$.
Many proofs of the isoperimetric inequality have been found. These employ a variety of techniques, including symmetrization techniques, variational methods, measure transportation (see [21], [26], [27], [28], [33]), and the Alexandrov-Bakelman-Pucci technique (see [10], [11], [12]).
In the measure transportation approach, it is convenient to normalize the domain $D$ so that $|D|=|B^{n}|$. The key idea is to construct a measure-preserving map $\Phi$ from the interior of $E$ to the open unit ball $B^{n}$ with the property that the eigenvalues of the differential of $\Phi$ are nonnegative real numbers. A map $\Phi$ with these properties can be constructed in several ways. Gromov’s work [21] uses the Knothe map, which has the additional property that the differential of $\Phi$ takes values in the set of upper triangular matrices. An alternative approach based on optimal transport was developed independently by McCann [26],[27] and Trudinger [33]. In this approach, one defines $\Phi(x)=\nabla u(x)$, where $u$ is a convex function solving the second boundary value problem for the Monge-Ampère equation. For this choice of $\Phi$, the differential of $\Phi$ takes values in the set of symmetric matrices.
In [11],[12], Cabré gave an elegant proof of the isoperimetric inequality based on what has become known as the Alexandrov-Bakelman-Pucci technique. The Alexandrov-Bakelman-Pucci technique in Cabré’s work is, in a certain sense, dual to the optimal transport approach developed by McCann and Trudinger. In the optimal transport approach, one considers a convex function satisfying $\det D^{2}u=1$ and concludes that $\Delta u\geq n$ by the arithmetic-geometric mean inequality. In the Alexandrov-Bakelman-Pucci approach, one considers a solution of $\Delta u=n$ and concludes that $\det D^{2}u\leq 1$ at each point where the Hessian of $u$ is weakly positive definite. The Alexandrov-Bakelman-Pucci technique only requires standard existence theorems for linear elliptic PDE.
In the remainder of this section, we describe a proof of Theorem 1.2 following the arguments in Cabré’s work [11],[12].
Without loss of generality, we may assume that $D$ is connected. By multiplying $f$ by a suitable constant, we may arrange that
(1) $\int_{D}|\nabla f|+\int_{\partial D}f=n\int_{D}f^{\frac{n}{n-1}}.$
We can find a function $u:D\to\mathbb{R}$ such that
(2) $\text{\rm div}(f\,\nabla u)=n\,f^{\frac{n}{n-1}}-|\nabla f|$
at each point in $D$ and
(3) $\langle\nabla u,\eta\rangle=1$
at each point on $\partial D$, where $\eta$ denotes the outward-pointing unit normal vector field to $\partial D$. The function $u$ is of class $C^{2,\gamma}$ for each $0<\gamma<1$. We define
$U$ $:=\{x\in D\setminus\partial D:|\nabla u(x)|<1\},$
$A$ $:=\{x\in U:D^{2}u(x)\geq 0\}.$
Moreover, we define a map $\Phi:U\to\mathbb{R}^{n}$ by $\Phi(x)=\nabla u(x)$ for all $x\in U$.
Lemma 1.3.
The open unit ball $B^{n}$ is contained in the image $\Phi(A)$.
Proof. Let us fix an arbitrary vector $\xi\in\mathbb{R}^{n}$ such that $|\xi|<1$. We define a function $w:D\to\mathbb{R}$ by $w(x):=u(x)-\langle x,\xi\rangle$. Using (3), we obtain
$\langle\nabla w(x),\eta(x)\rangle=\langle\nabla u(x),\eta(x)\rangle-\langle\eta(x),\xi\rangle=1-\langle\eta(x),\xi\rangle>0$
for each point $x\in\partial D$. Consequently, we can find a point $\bar{x}\in D\setminus\partial D$ where the function $w$ attains its minimum. Clearly, $\nabla w(\bar{x})=0$ and $D^{2}w(\bar{x})\geq 0$. From this, we deduce that $\xi=\nabla u(\bar{x})$ and $D^{2}u(\bar{x})\geq 0$. This implies $\bar{x}\in A$ and $\Phi(\bar{x})=\xi$. Thus, $B^{n}\subset\Phi(A)$.
Lemma 1.4.
The Jacobian determinant of $\Phi$ satisfies
$0\leq\det D\Phi(x)\leq f(x)^{\frac{n}{n-1}}$
for each point $x\in A$.
Proof. Let us fix a point $x\in A$. Using (2), we obtain
$\Delta u(x)$ $=n\,f(x)^{\frac{1}{n-1}}-f(x)^{-1}\,|\nabla f(x)|-f(x)^{-1}\,\langle\nabla f(x),\nabla u(x)\rangle$
$\leq n\,f(x)^{\frac{1}{n-1}}.$
Moreover, $D^{2}u(x)\geq 0$ since $x\in A$. Hence, the arithmetic-geometric mean inequality implies
$0\leq\det D^{2}u(x)\leq\Big(\frac{\Delta u(x)}{n}\Big)^{n}\leq f(x)^{\frac{n}{n-1}}.$
This completes the proof of Lemma 1.4.
Using Lemma 1.3 and Lemma 1.4, we obtain
$|B^{n}|\leq\int_{A}|\det D\Phi(x)|\,dx\leq\int_{D}f(x)^{\frac{n}{n-1}}\,dx.$
Using this estimate together with the normalization condition (1), we conclude that
$\int_{D}|\nabla f|+\int_{\partial D}f=n\int_{D}f^{\frac{n}{n-1}}\geq n\,|B^{n}|^{\frac{1}{n}}\,\Big(\int_{D}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}}.$
This completes the proof of Theorem 1.2.
2. The Fenchel-Willmore-Chen inequality for submanifolds of Euclidean space
In this section, we discuss how the well-known Fenchel-Willmore-Chen inequality fits within the framework of the Alexandrov-Bakelman-Pucci technique.
Theorem 2.1 (Fenchel-Willmore-Chen inequality).
Let $\Sigma$ be a compact $n$-dimensional submanifold of $\mathbb{R}^{n+m}$ without boundary. Then
$\int_{\Sigma}\Big(\frac{|H|}{n}\Big)^{n}\geq|S^{n}|,$
where $H$ denotes the mean curvature vector of $\Sigma$.
Theorem 2.1 was proved by Chen [15], building upon earlier work of Fenchel [20] and Willmore [34].
In the remainder of this section, we describe the proof of Theorem 2.1. We define
$U$ $:=\{(x,y):x\in\Sigma,\,y\in T_{x}^{\perp}\Sigma,\,|y|<1\},$
$A$ $:=\{(x,y)\in U:-\langle I\!I(x),y\rangle\geq 0\}.$
Moreover, we define a map $\Phi:U\to\mathbb{R}^{n+m}$ by $\Phi(x,y)=y$ for all $(x,y)\in U$.
Lemma 2.2.
The open unit ball $B^{n+m}$ is contained in the image $\Phi(A)$.
Proof. Let us fix an arbitrary vector $\xi\in\mathbb{R}^{n+m}$ such that $|\xi|<1$. We define a function $w:\Sigma\to\mathbb{R}$ by $w(x):=-\langle x,\xi\rangle$. We can find a point $\bar{x}\in\Sigma$ where the function $w$ attains its minimum. Clearly, $\nabla^{\Sigma}w(\bar{x})=0$ and $D_{\Sigma}^{2}w(\bar{x})\geq 0$. From this, we deduce that $\xi$ is orthogonal to the tangent space $T_{\bar{x}}\Sigma$ and $\langle I\!I(\bar{x}),\xi\rangle\geq 0$. Therefore, if we define a normal vector $\bar{y}\in T_{\bar{x}}^{\perp}\Sigma$ by $\bar{y}=\xi$, then $(\bar{x},\bar{y})\in A$ and $\Phi(\bar{x},\bar{y})=\xi$. Thus, $B^{n+m}\subset\Phi(A)$.
Lemma 2.3.
The Jacobian determinant of $\Phi$ is given by
$\det D\Phi(x,y)=\det(-\langle I\!I(x),y\rangle)$
for each point $(x,y)\in U$.
Proof. Let us fix a point $(\bar{x},\bar{y})\in U$. Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal basis of the tangent space $T_{\bar{x}}\Sigma$, and let $(x_{1},\ldots,x_{n})$ be a local coordinate system on $\Sigma$ such that $\frac{\partial}{\partial x_{i}}=e_{i}$ at the point $\bar{x}$. Moreover, let $\{\nu_{1},\ldots,\nu_{m}\}$ denote a local orthonormal frame for the normal bundle $T^{\perp}\Sigma$. Every normal vector $y$ can be written in the form $y=\sum_{\alpha=1}^{m}y_{\alpha}\nu_{\alpha}$. With this understood, $(x_{1},\ldots,x_{n},y_{1},\ldots,y_{m})$ is a local coordinate system on the total space of the normal bundle $T^{\perp}\Sigma$. A straightforward calculation gives
$\Big\langle\frac{\partial\Phi}{\partial x_{i}}(\bar{x},\bar{y}),e_{j}\Big\rangle$ $=-\langle I\!I(e_{i},e_{j}),\bar{y}\rangle,$
$\Big\langle\frac{\partial\Phi}{\partial y_{\alpha}}(\bar{x},\bar{y}),e_{j}\Big\rangle$ $=0,$
$\Big\langle\frac{\partial\Phi}{\partial y_{\alpha}}(\bar{x},\bar{y}),\nu_{\beta}\Big\rangle$ $=\delta_{\alpha\beta}.$
In other words, the differential of $\Phi$ at the point $(\bar{x},\bar{y})$ takes the form
$D\Phi(\bar{x},\bar{y})=\begin{bmatrix}-\langle I\!I(\bar{x}),\bar{y}\rangle&0\\ *&\text{\rm id}\end{bmatrix}$.
The assertion follows by taking the determinant.
Lemma 2.4.
The Jacobian determinant of $\Phi$ satisfies
$0\leq\det D\Phi(x,y)\leq\Big(-\frac{\langle H(x),y\rangle}{n}\Big)^{n}$
for each point $(x,y)\in A$.
Proof. Let us fix a point $(x,y)\in A$. Then $\langle I\!I(x),y\rangle\geq 0$. Hence, the arithmetic-geometric mean inequality implies
$0\leq\det(-\langle I\!I(x),y\rangle)\leq\Big(-\frac{\langle H(x),y\rangle}{n}\Big)^{n}.$
The assertion follows now from Lemma 2.3.
Lemma 2.5.
We have
$\int_{\{y\in\mathbb{R}^{m}:|y|<1\}}(-\langle a,y\rangle)_{+}^{n}\,dy=\frac{|B^{n+m}|}{|S^{n}|}\,|a|^{n}$
for each $a\in\mathbb{R}^{m}$.
Proof. It suffices to prove the assertion in the special case when $a\in\mathbb{R}^{m}$ is a unit vector. For each unit vector $a\in\mathbb{R}^{m}$, we obtain
$\int_{\{y\in\mathbb{R}^{m}:|y|<1\}}(-\langle a,y\rangle)_{+}^{n}\,dy$ $=|B^{m-1}|\int_{0}^{1}s^{n}\,(1-s^{2})^{\frac{m-1}{2}}\,ds$
$=\frac{1}{2}\,|B^{m-1}|\int_{0}^{1}t^{\frac{n-1}{2}}\,(1-t)^{\frac{m-1}{2}}\,dt$
$=\frac{\Gamma(\frac{n+1}{2})\,\Gamma(\frac{m+1}{2})}{2\,\Gamma(\frac{n+m+2}{2})}\,|B^{m-1}|.$
The first equality follows from Fubini’s theorem. In the second equality, we have used the substitution $t=s^{2}$. The third equality follows from the well-known formula relating the Beta function to the Gamma function. Using the identities
$|S^{n}|=\frac{2\,\pi^{\frac{n+1}{2}}}{\Gamma(\frac{n+1}{2})},$
$|B^{m-1}|=\frac{\pi^{\frac{m-1}{2}}}{\Gamma(\frac{m+1}{2})},$
and
$|B^{n+m}|=\frac{\pi^{\frac{n+m}{2}}}{\Gamma(\frac{n+m+2}{2})},$
we obtain
$\frac{\Gamma(\frac{n+1}{2})\,\Gamma(\frac{m+1}{2})}{2\,\Gamma(\frac{n+m+2}{2})}\,|B^{m-1}|=\frac{|B^{n+m}|}{|S^{n}|}.$
This completes the proof of Lemma 2.5.
Using Lemma 2.2 and Lemma 2.4, we obtain
$|B^{n+m}|$ $\leq\int_{\Sigma}\bigg(\int_{T_{x}^{\perp}\Sigma}|\det D\Phi(x,y)|\,1_{A}(x,y)\,dy\bigg)\,d\text{\rm vol}(x)$
$\leq\int_{\Sigma}\bigg(\int_{\{y\in T_{x}^{\perp}\Sigma:|y|<1\}}\Big(-\frac{\langle H(x),y\rangle}{n}\Big)_{+}^{n}\,dy\bigg)\,d\text{\rm vol}(x).$
On the other hand, Lemma 2.5 implies that
$\int_{\{y\in T_{x}^{\perp}\Sigma:|y|<1\}}(-\langle H(x),y\rangle)_{+}^{n}\,dy=\frac{|B^{n+m}|}{|S^{n}|}\,|H(x)|^{n}$
for each point $x\in\Sigma$. Putting these facts together, we conclude that
$|B^{n+m}|\leq\frac{|B^{n+m}|}{|S^{n}|}\int_{\Sigma}\Big(\frac{|H(x)|}{n}\Big)^{n}\,d\text{\rm vol}(x).$
This completes the proof of Theorem 2.1.
3. The Sobolev inequality for submanifolds of Euclidean space
In this section, we explain how the Alexandrov-Bakelman-Pucci technique can be used to prove a isoperimetric inequality for submanifolds.
Theorem 3.1 (cf. [5]).
Let $\Sigma$ be a compact $n$-dimensional submanifold of $\mathbb{R}^{n+m}$ (possibly with boundary), where $m\geq 2$. Let $f$ be a smooth positive function on $\Sigma$. Then
$\int_{\Sigma}\sqrt{|\nabla^{\Sigma}f|^{2}+f^{2}\,|H|^{2}}+\int_{\Sigma}f\geq n\,\Big(\frac{(n+m)\,|B^{n+m}|}{m\,|B^{m}|}\Big)^{\frac{1}{n}}\,\Big(\int_{\Sigma}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}},$
where $H$ denotes the mean curvature vector of $\Sigma$.
The case $m=2$ is of particular interest. Using Theorem 3.1 and the formula $(n+2)\,|B^{n+2}|=2\pi\,|B^{n}|=2\,|B^{2}|\,|B^{n}|$, we obtain a sharp Sobolev inequality for submanifolds of codimension $2$.
Corollary 3.2 (cf. [5]).
Let $\Sigma$ be a compact $n$-dimensional submanifold of $\mathbb{R}^{n+2}$ (possibly with boundary). Let $f$ be a smooth positive function on $\Sigma$. Then
$\int_{\Sigma}\sqrt{|\nabla^{\Sigma}f|^{2}+f^{2}\,|H|^{2}}+\int_{\Sigma}f\geq n\,|B^{n}|^{\frac{1}{n}}\,\Big(\int_{\Sigma}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}},$
where $H$ denotes the mean curvature vector of $\Sigma$.
In the special case when $\Sigma$ is a minimal submanifold, we can draw the following conclusion.
Corollary 3.3 (cf. [5]).
Let $\Sigma$ be a compact $n$-dimensional minimal submanifold of $\mathbb{R}^{n+2}$ (possibly with boundary). Then $|\partial\Sigma|\geq n\,|B^{n}|^{\frac{1}{n}}\,|\Sigma|^{\frac{n-1}{n}}$.
Corollary 3.2 and Corollary 3.3 immediately imply the corresponding inequalities for submanifolds of codimension $1$. A direct proof in the codimension $1$ setting can be found in [8].
The isoperimetric inequality for minimal surfaces has a long history, going back to the seminal work of Carleman [13] in 1921. A number of partial results have been established over the years; see [3], [16], [19], [23], [24], [30], [31], [32]. In particular, Almgren [3] showed that the sharp isoperimetric inequality holds for area-minimizing submanifolds of arbitrary dimension and codimension. Almgren’s proof relies on arguments from geometric measure theory together with a generalization of the Fenchel-Willmore-Chen inequality.
Corollary 3.2 can be viewed as a sharp version of the famous Michael-Simon Sobolev inequality. The non-sharp version of the inequality was proved independently by Allard [2] and by Michael and Simon [29] (see also [14]).
In the remainder of this section, we describe a proof of Theorem 3.1 using the Alexandrov-Bakelman-Pucci technique, following the arguments in [5]. We refer to [9] for an alternative approach based on an optimal transport problem between spaces of unequal dimension.
Without loss of generality, we assume that $\Sigma$ is connected. By multiplying $f$ by a suitable constant, we may arrange that
(4) $\int_{\Sigma}\sqrt{|\nabla^{\Sigma}f|^{2}+f^{2}\,|H|^{2}}+\int_{\partial\Sigma}f=n\int_{\Sigma}f^{\frac{n}{n-1}}.$
We can find a function $u:\Sigma\to\mathbb{R}$ with the property that
(5) $\text{\rm div}_{\Sigma}(f\,\nabla^{\Sigma}u)=n\,f^{\frac{n}{n-1}}-\sqrt{|\nabla^{\Sigma}f|^{2}+f^{2}\,|H|^{2}}$
at each point on $\Sigma$ and
(6) $\langle\nabla^{\Sigma}u,\eta\rangle=1$
at each point on $\partial\Sigma$, where $\eta$ denotes the outward-pointing unit normal vector field to $\partial\Sigma$ in $\Sigma$. The function $u$ is of class $C^{2,\gamma}$ for each $0<\gamma<1$. We define
$\Omega$ $:=\{x\in\Sigma\setminus\partial\Sigma:|\nabla^{\Sigma}u(x)|<1\},$
$U$ $:=\{(x,y):x\in\Sigma\setminus\partial\Sigma,\,y\in T_{x}^{\perp}\Sigma,\,|\nabla^{\Sigma}u(x)|^{2}+|y|^{2}<1\},$
$A$ $:=\{(x,y)\in U:D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle\geq 0\}.$
Moreover, we define a map $\Phi:U\to\mathbb{R}^{n+m}$ by $\Phi(x,y)=\nabla^{\Sigma}u(x)+y$ for all $(x,y)\in U$. Note that $|\Phi(x,y)|^{2}=|\nabla^{\Sigma}u(x)|^{2}+|y|^{2}<1$ for all $(x,y)\in U$.
Lemma 3.4.
The open unit ball $B^{n+m}$ is contained in the image $\Phi(A)$.
Proof. Let us fix an arbitrary vector $\xi\in\mathbb{R}^{n+m}$ such that $|\xi|<1$. We define a function $w:\Sigma\to\mathbb{R}$ by $w(x):=u(x)-\langle x,\xi\rangle$. Using (6), we obtain
$\langle\nabla^{\Sigma}w(x),\eta(x)\rangle=\langle\nabla^{\Sigma}u(x),\eta(x)\rangle-\langle\eta(x),\xi\rangle=1-\langle\eta(x),\xi\rangle>0$
for each point $x\in\partial\Sigma$. Consequently, we can find a point $\bar{x}\in\Sigma\setminus\partial\Sigma$ where the function $w$ attains its minimum. Clearly, $\nabla^{\Sigma}w(\bar{x})=0$ and $D_{\Sigma}^{2}w(\bar{x})\geq 0$. Therefore, we can find a vector $\bar{y}\in T_{\bar{x}}^{\perp}\Sigma$ such that $\xi=\nabla^{\Sigma}u(\bar{x})+\bar{y}$ and $D_{\Sigma}^{2}u(\bar{x})-\langle I\!I(\bar{x}),\bar{y}\rangle\geq 0$. This implies $(\bar{x},\bar{y})\in A$ and $\Phi(\bar{x},\bar{y})=\xi$. Thus, $B^{n+m}\subset\Phi(A)$.
Lemma 3.5.
The Jacobian determinant of $\Phi$ is given by
$\det D\Phi(x,y)=\det(D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle)$
for all $(x,y)\in U$.
Proof. Let us fix a point $(\bar{x},\bar{y})\in U$. Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal basis of the tangent space $T_{\bar{x}}\Sigma$, and let $(x_{1},\ldots,x_{n})$ be a local coordinate system on $\Sigma$ such that $\frac{\partial}{\partial x_{i}}=e_{i}$ at the point $\bar{x}$. Moreover, let $\{\nu_{1},\ldots,\nu_{m}\}$ denote a local orthonormal frame for the normal bundle $T^{\perp}\Sigma$. Every normal vector $y$ can be written in the form $y=\sum_{\alpha=1}^{m}y_{\alpha}\nu_{\alpha}$. With this understood, $(x_{1},\ldots,x_{n},y_{1},\ldots,y_{m})$ is a local coordinate system on the total space of the normal bundle $T^{\perp}\Sigma$. A straightforward calculation gives
$\Big\langle\frac{\partial\Phi}{\partial x_{i}}(\bar{x},\bar{y}),e_{j}\Big\rangle$ $=(D_{\Sigma}^{2}u)(e_{i},e_{j})-\langle I\!I(e_{i},e_{j}),\bar{y}\rangle,$
$\Big\langle\frac{\partial\Phi}{\partial y_{\alpha}}(\bar{x},\bar{y}),e_{j}\Big\rangle$ $=0,$
$\Big\langle\frac{\partial\Phi}{\partial y_{\alpha}}(\bar{x},\bar{y}),\nu_{\beta}\Big\rangle$ $=\delta_{\alpha\beta}.$
In other words, the differential of $\Phi$ at the point $(\bar{x},\bar{y})$ takes the form
$D\Phi(\bar{x},\bar{y})=\begin{bmatrix}D_{\Sigma}^{2}u(\bar{x})-\langle I\!I(\bar{x}),\bar{y}\rangle&0\\ *&\text{\rm id}\end{bmatrix}$.
The assertion follows by taking the determinant.
Lemma 3.6.
The Jacobian determinant of $\Phi$ satisfies
$0\leq\det D\Phi(x,y)\leq f(x)^{\frac{n}{n-1}}$
for all $(x,y)\in A$.
Proof. Let us fix a point $(x,y)\in A$. Then $|\nabla^{\Sigma}u(x)|^{2}+|y|^{2}<1$. Using the Cauchy-Schwarz inequality, we obtain
$-\langle\nabla^{\Sigma}f(x),\nabla^{\Sigma}u(x)\rangle-f(x)\,\langle H(x),y\rangle$
$\leq\sqrt{|\nabla^{\Sigma}f(x)|^{2}+f(x)^{2}\,|H(x)|^{2}}\,\sqrt{|\nabla^{\Sigma}u(x)|^{2}+|y|^{2}}$
$\leq\sqrt{|\nabla^{\Sigma}f(x)|^{2}+f(x)^{2}\,|H(x)|^{2}}.$
Using (5), we obtain
$\Delta_{\Sigma}u(x)-\langle H(x),y\rangle$
$=n\,f(x)^{\frac{1}{n-1}}-f(x)^{-1}\,\sqrt{|\nabla^{\Sigma}f(x)|^{2}+f(x)^{2}\,|H(x)|^{2}}$
$-f(x)^{-1}\,\langle\nabla^{\Sigma}f(x),\nabla^{\Sigma}u(x)\rangle-\langle H(x),y\rangle$
$\leq n\,f(x)^{\frac{1}{n-1}}.$
Moreover, $D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle\geq 0$ since $(x,y)\in A$. Hence, the arithmetic-geometric mean inequality implies
$0\leq\det(D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle)\leq\Big(\frac{\text{\rm tr}(D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle)}{n}\Big)^{n}\leq f(x)^{\frac{n}{n-1}}.$
The assertion follows now from Lemma 3.5. This completes the proof of Lemma 3.6.
Using Lemma 3.4 and Lemma 3.6, we obtain
$|B^{n+m}|\,(1-\sigma^{n+m})$
$=\int_{\{\xi\in\mathbb{R}^{n+m}:\sigma^{2}<|\xi|^{2}<1\}}1\,d\xi$
$\leq\int_{\Omega}\bigg(\int_{\{y\in T_{x}^{\perp}\Sigma:\sigma^{2}<|\Phi(x,y)|^{2}<1\}}|\det D\Phi(x,y)|\,1_{A}(x,y)\,dy\bigg)\,d\text{\rm vol}(x)$
$\leq\int_{\Omega}\bigg(\int_{\{y\in T_{x}^{\perp}\Sigma:\sigma^{2}<|\nabla^{\Sigma}u(x)|^{2}+|y|^{2}<1\}}f(x)^{\frac{n}{n-1}}\,dy\bigg)\,d\text{\rm vol}(x)$
$=|B^{m}|\int_{\Omega}\Big[(1-|\nabla^{\Sigma}u(x)|^{2})^{\frac{m}{2}}-(\sigma^{2}-|\nabla^{\Sigma}u(x)|^{2})_{+}^{\frac{m}{2}}\Big]\,f(x)^{\frac{n}{n-1}}\,d\text{\rm vol}(x)$
for all $0\leq\sigma<1$. Since $m\geq 2$, an elementary inequality gives
$(1-|\nabla^{\Sigma}u(x)|^{2})^{\frac{m}{2}}-(\sigma^{2}-|\nabla^{\Sigma}u(x)|^{2})_{+}^{\frac{m}{2}}$
$\leq\frac{m}{2}\,\Big[(1-|\nabla^{\Sigma}u(x)|^{2})-(\sigma^{2}-|\nabla^{\Sigma}u(x)|^{2})_{+}\Big]$
$\leq\frac{m}{2}\,(1-\sigma^{2})$
for all $x\in\Omega$ and all $0\leq\sigma<1$. Consequently,
$|B^{n+m}|\,(1-\sigma^{n+m})\leq\frac{m}{2}\,|B^{m}|\,(1-\sigma^{2})\int_{\Sigma}f^{\frac{n}{n-1}}$
for all $0\leq\sigma<1$. Finally, we divide by $1-\sigma$ and take the limit as $\sigma\to 1$. Thus,
$(n+m)\,|B^{n+m}|\leq m\,|B^{m}|\int_{\Sigma}f^{\frac{n}{n-1}}.$
Using this estimate together with the normalization condition (4), we conclude that
$\int_{\Sigma}\sqrt{|\nabla^{\Sigma}f|^{2}+f^{2}\,|H|^{2}}+\int_{\partial\Sigma}f$
$=n\int_{\Sigma}f^{\frac{n}{n-1}}\geq n\,\Big(\frac{(n+m)\,|B^{n+m}|}{m\,|B^{m}|}\Big)^{\frac{1}{n}}\,\Big(\int_{\Sigma}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}}.$
This completes the proof of Theorem 3.1.
4. The logarithmic Sobolev inequality for submanifolds of Euclidean space
In this section, we discuss a sharp logarithmic Sobolev inequality for submanifolds.
Theorem 4.1 (cf. [6]).
Let $\Sigma$ be a compact $n$-dimensional submanifold of $\mathbb{R}^{n+m}$ without boundary. Let $f$ be a positive smooth function on $\Sigma$. Then
$\int_{\Sigma}f\,\Big(\log f+n+\frac{n}{2}\,\log(4\pi)\Big)-\int_{\Sigma}f^{-1}\,|\nabla^{\Sigma}f|^{2}-\int_{\Sigma}f\,|H|^{2}$
$\leq\Big(\int_{\Sigma}f\Big)\,\log\Big(\int_{\Sigma}f\Big),$
where $H$ denotes the mean curvature vector of $\Sigma$.
If we put $f=(4\pi)^{-\frac{n}{2}}\,e^{-\frac{1}{4}\,|x|^{2}}\,\varphi$, then Theorem 4.1 can be rewritten in the following equivalent form.
Corollary 4.2 (cf. [6]).
Let $\Sigma$ be a compact $n$-dimensional submanifold of $\mathbb{R}^{n+m}$ without boundary, and let
$d\gamma=(4\pi)^{-\frac{n}{2}}\,e^{-\frac{1}{4}\,|x|^{2}}\,d\text{\rm vol}$
denote the Gaussian measure on $\Sigma$. Let $\varphi$ be a positive smooth function on $\Sigma$. Then
$\int_{\Sigma}\varphi\,\log\varphi\,d\gamma-\int_{\Sigma}\varphi^{-1}\,|\nabla^{\Sigma}\varphi|^{2}\,d\gamma-\int_{\Sigma}\varphi\,\Big|H+\frac{1}{2}\,x^{\perp}\Big|^{2}\,d\gamma$
$\leq\Big(\int_{\Sigma}\varphi\,d\gamma\Big)\,\log\Big(\int_{\Sigma}\varphi\,d\gamma\Big),$
where $H$ denotes the mean curvature vector of $\Sigma$.
Corollary 4.2 can be viewed as a sharp version of Ecker’s logarithmic Sobolev inequality for submanifolds [18].
In the remainder of this section, we explain how Theorem 4.1 can be proved using the Alexandrov-Bakelman-Pucci technique.
Without loss of generality, we may assume that $\Sigma$ is connected. By multiplying $f$ by a suitable constant, we may arrange that
(7) $\int_{\Sigma}f\,\log f-\int_{\Sigma}f^{-1}\,|\nabla^{\Sigma}f|^{2}-\int_{\Sigma}f\,|H|^{2}=0.$
We can find a smooth function $u:\Sigma\to\mathbb{R}$ such that
(8) $\text{\rm div}_{\Sigma}(f\,\nabla^{\Sigma}u)=f\,\log f-f^{-1}\,|\nabla^{\Sigma}f|^{2}-f\,|H|^{2}.$
We define
$U$ $:=\{(x,y):x\in\Sigma,\,y\in T_{x}^{\perp}\Sigma\},$
$A$ $:=\{(x,y)\in U:D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle\geq 0\}.$
Moreover, we define a map $\Phi:U\to\mathbb{R}^{n+m}$ by $\Phi(x,y)=\nabla^{\Sigma}u(x)+y$ for all $(x,y)\in U$. Note that $|\Phi(x,y)|^{2}=|\nabla^{\Sigma}u(x)|^{2}+|y|^{2}$ for all $(x,y)\in U$. Arguing as in the proof of Lemma 3.4, we can show that $\Phi(A)=\mathbb{R}^{n+m}$. Moreover, the Jacobian determinant of $\Phi$ is given by
(9) $\det D\Phi(x,y)=\det(D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle)$
for each point $(x,y)\in U$.
Lemma 4.3.
The Jacobian determinant of $\Phi$ satisfies
$0\leq e^{-\frac{1}{4}\,|\Phi(x,y)|^{2}}\,\det D\Phi(x,y)\leq f(x)\,e^{-\frac{1}{4}\,|2H(x)+y|^{2}-n}$
for each point $(x,y)\in A$.
Proof. Let us fix a point $(x,y)\in A$. Using (8), we obtain
$\Delta_{\Sigma}u(x)-\langle H(x),y\rangle$
$=\log f(x)-f(x)^{-2}\,|\nabla^{\Sigma}f(x)|^{2}-|H(x)|^{2}$
$-f(x)^{-1}\,\langle\nabla^{\Sigma}f(x),\nabla^{\Sigma}u(x)\rangle-\langle H(x),y\rangle$
$=\log f(x)+\frac{1}{4}\,(|\nabla^{\Sigma}u(x)|^{2}+|y|^{2})$
$-\frac{1}{4}\,f(x)^{-2}\,|2\,\nabla^{\Sigma}f(x)+f(x)\,\nabla^{\Sigma}u(x)|^{2}-\frac{1}{4}\,|2H(x)+y|^{2}$
$\leq\log f(x)+\frac{1}{4}\,(|\nabla^{\Sigma}u(x)|^{2}+|y|^{2})-\frac{1}{4}\,|2H(x)+y|^{2}.$
Moreover, $D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle\geq 0$ since $(x,y)\in A$. Using the elementary inequality $\lambda\leq e^{\lambda-1}$, we obtain
$0$ $\leq\det(D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle)$
$\leq e^{\text{\rm tr}(D_{\Sigma}^{2}u(x)-\langle I\!I(x),y\rangle)-n}$
$\leq f(x)\,e^{\frac{1}{4}\,(|\nabla^{\Sigma}u(x)|^{2}+|y|^{2})-\frac{1}{4}\,|2H(x)+y|^{2}-n}.$
Using (9), we conclude that
$0\leq\det D\Phi(x,y)\leq f(x)\,e^{\frac{1}{4}\,|\Phi(x,y)|^{2}-\frac{1}{4}\,|2H(x)+y|^{2}-n}.$
This completes the proof of Lemma 4.3.
Using Lemma 4.3 and the fact that $\Phi(A)=\mathbb{R}^{n+m}$, we obtain
$1$ $=(4\pi)^{-\frac{n+m}{2}}\int_{\mathbb{R}^{n+m}}e^{-\frac{1}{4}\,|\xi|^{2}}\,d\xi$
$\leq(4\pi)^{-\frac{n+m}{2}}\int_{\Sigma}\bigg(\int_{T_{x}^{\perp}\Sigma}e^{-\frac{1}{4}\,|\Phi(x,y)|^{2}}\,|\det D\Phi(x,y)|\,1_{A}(x,y)\,dy\bigg)\,d\text{\rm vol}(x)$
$\leq(4\pi)^{-\frac{n+m}{2}}\int_{\Sigma}\bigg(\int_{T_{x}^{\perp}\Sigma}f(x)\,e^{-\frac{1}{4}\,|2H(x)+y|^{2}-n}\,dy\bigg)\,d\text{\rm vol}(x)$
$=(4\pi)^{-\frac{n}{2}}\,e^{-n}\int_{\Sigma}f(x)\,d\text{\rm vol}(x).$
Rearranging terms gives
$n+\frac{n}{2}\,\log(4\pi)\leq\log\Big(\int_{\Sigma}f\Big).$
Using this estimate together with the normalization condition (7), we conclude that
$\int_{\Sigma}f\,\Big(\log f+n+\frac{n}{2}\,\log(4\pi)\Big)-\int_{\Sigma}f^{-1}\,|\nabla^{\Sigma}f|^{2}-\int_{\Sigma}f\,|H|^{2}$
$=\int_{\Sigma}f\,\Big(n+\frac{n}{2}\,\log(4\pi)\Big)$
$\leq\Big(\int_{\Sigma}f\Big)\,\log\Big(\int_{\Sigma}f\Big).$
This completes the proof of Theorem 4.1.
5. The Sobolev inequality for manifolds with nonnegative Ricci curvature
In this section, we discuss how the isoperimetric inequality in Euclidean space can be generalized to manifolds with nonnegative Ricci curvature. Let $(M,g)$ be a complete noncompact manifold of dimension $n$ with nonnegative Ricci curvature, and let $q$ be an arbitrary point in $M$. It follows from the Bishop-Gromov relative volume comparison theorem that the function
$r\mapsto\frac{|\{p\in M:d(p,q)<r\}|}{r^{n}}$
is monotone decreasing. In particular, we can find a real number $\theta\in[0,1]$ such that
$\frac{|\{p\in M:d(p,q)<r\}|}{r^{n}}\to|B^{n}|\,\theta$
as $r\to\infty$. It is easy to see that $\theta$ is independent of the choice of the point $q$. We refer to $\theta$ as the asymptotic volume ratio of $(M,g)$.
Theorem 5.1 (cf. [1], [7], [17]).
Let $(M,g)$ be a complete noncompact manifold of dimension $n$ with nonnegative Ricci curvature. Let $D$ be a compact domain in $M$ with boundary $\partial D$. Then $|\partial D|\geq n\,|B^{n}|^{\frac{1}{n}}\,\theta^{\frac{1}{n}}\,|D|^{\frac{n-1}{n}}$, where $\theta$ denotes the asymptotic volume ratio of $(M,g)$.
In the three-dimensional case, Theorem 5.1 was proved by Agostiniani, Fogagnolo, and Mazzieri (cf. [1], Theorem 6.1). Their proof builds on an argument due to Huisken and uses geometric flows. In [7], we gave a proof of Theorem 5.1 in all dimensions based on the Alexandrov-Bakelman-Pucci technique. Finally, Balogh and Kristály [4] observed that Theorem 5.1 can alternatively be deduced from the Riemannian interpolation inequality of Cordero-Erausquin, McCann, and Schmuckenschläger [17].
Theorem 5.1 is a special case of the following Sobolev inequality.
Theorem 5.2.
Let $(M,g)$ be a complete noncompact manifold of dimension $n$ with nonnegative Ricci curvature. Let $D$ be a compact domain in $M$ with boundary $\partial D$, and let $f$ be a positive smooth function on $D$. Then
$\int_{D}|\nabla f|+\int_{\partial D}f\geq n\,|B^{n}|^{\frac{1}{n}}\,\theta^{\frac{1}{n}}\,\Big(\int_{D}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}},$
where $\theta$ denotes the asymptotic volume ratio of $(M,g)$.
In the remainder of this section, we describe the proof of Theorem 5.2 based on the Alexandrov-Bakelman-Pucci maximum principle. We will follow the arguments in [7].
Without loss of generality, we may assume that $D$ is connected. By multiplying $f$ by a suitable constant, we may arrange that
(10) $\int_{D}|\nabla f|+\int_{\partial D}f=n\int_{D}f^{\frac{n}{n-1}}.$
We can find a function $u:D\to\mathbb{R}$ such that
(11) $\text{\rm div}(f\,\nabla u)=n\,f^{\frac{n}{n-1}}-|\nabla f|$
at each point in $D$ and
(12) $\langle\nabla u,\eta\rangle=1$
at each point on $\partial D$, where $\eta$ denotes the outward-pointing unit normal to $\partial D$. The function $u$ is of class $C^{2,\gamma}$ for each $0<\gamma<1$. We define
$U:=\{x\in D\setminus\partial D:|\nabla u(x)|<1\}.$
For each $r>0$, we denote by $A_{r}$ the set of all points $\bar{x}\in U$ with the property that
$r\,u(x)+\frac{1}{2}\,d\big(x,\exp_{\bar{x}}(r\,\nabla u(\bar{x}))\big)^{2}\geq r\,u(\bar{x})+\frac{1}{2}\,r^{2}\,|\nabla u(\bar{x})|^{2}$
for all $x\in D$. Moreover, for each $r>0$, we define a map $\Phi_{r}:U\to M$ by
$\Phi_{r}(x)=\exp_{x}(r\,\nabla u(x))$
for all $x\in U$.
Lemma 5.3.
Assume that $x\in U$. Then $\Delta u(x)\leq n\,f(x)^{\frac{1}{n-1}}$.
Proof. Let us fix a point $x\in U$. Using (2), we obtain
$\Delta u(x)$ $=n\,f(x)^{\frac{1}{n-1}}-f(x)^{-1}\,|\nabla f(x)|-f(x)^{-1}\,\langle\nabla f(x),\nabla u(x)\rangle$
$\leq n\,f(x)^{\frac{1}{n-1}}.$
This completes the proof of Lemma 5.3.
Lemma 5.4.
For each $r>0$, the set
$\Big\{p\in M:\sup_{x\in D}d(x,p)<r\Big\}$
is contained in the image $\Phi_{r}(A_{r})$.
Proof. Let us fix a real number $r>0$. Let $p$ be a point in $M$ with the property that $\sup_{x\in D}d(x,p)<r$. We can find a point $\bar{x}\in D$ such that
$r\,u(\bar{x})+\frac{1}{2}\,d(\bar{x},p)^{2}=\inf_{x\in D}(r\,u(x)+\frac{1}{2}\,d(x,p)^{2}).$
Using (12) and the inequality $\sup_{x\in D}d(x,p)<r$, we conclude that the point $\bar{x}$ lies in the interior of $D$. Let $\bar{\gamma}:[0,r]\to M$ be a minimizing geodesic such that $\bar{\gamma}(0)=\bar{x}$ and $\bar{\gamma}(r)=p$. The path $\bar{\gamma}$ minimizes the functional $u(\gamma(0))+\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=p$. The first variation formula implies that
$\nabla u(\bar{x})=\bar{\gamma}^{\prime}(0).$
Consequently,
$\exp_{\bar{x}}(r\,\nabla u(\bar{x}))=\exp_{\bar{\gamma}(0)}(r\,\bar{\gamma}^{\prime}(0))=\bar{\gamma}(r)=p.$
Moreover,
$r\,|\nabla u(\bar{x})|=r\,|\bar{\gamma}^{\prime}(0)|=d(\bar{x},p).$
Since $d(\bar{x},p)<r$, it follows that $|\nabla u(\bar{x})|<1$. Therefore, $\bar{x}\in U$. Finally, for each point $x\in D$, we have
$r\,u(x)+\frac{1}{2}\,d\big(x,\exp_{\bar{x}}(r\,\nabla u(\bar{x}))\big)^{2}$ $=r\,u(x)+\frac{1}{2}\,d(x,p)^{2}$
$\geq r\,u(\bar{x})+\frac{1}{2}\,d(\bar{x},p)^{2}$
$=r\,u(\bar{x})+\frac{1}{2}\,r^{2}\,|\nabla u(\bar{x})|^{2}.$
Thus, $\bar{x}\in A_{r}$ and $\Phi_{r}(\bar{x})=p$. This completes the proof of Lemma 5.4.
Lemma 5.5.
Assume that $r>0$ and $\bar{x}\in A_{r}$. We define $\bar{\gamma}(t):=\exp_{\bar{x}}(t\,\nabla u(\bar{x}))$ for all $t\in[0,r]$. If $Z$ is a smooth vector field along $\bar{\gamma}$ satisfying $Z(r)=0$, then
$(D^{2}u)(Z(0),Z(0))+\int_{0}^{r}\big(|D_{t}Z(t)|^{2}-R(\bar{\gamma}^{\prime}(t),Z(t),\bar{\gamma}^{\prime}(t),Z(t))\big)\,dt\geq 0.$
Proof. Since $\bar{x}\in A_{r}$, the path $\bar{\gamma}$ minimizes the functional $u(\gamma(0))+\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=\bar{\gamma}(r)$. Hence, the assertion follows from the second variation formula.
Lemma 5.6.
Assume that $r>0$ and $\bar{x}\in A_{r}$. We define $\bar{\gamma}(t):=\exp_{\bar{x}}(t\,\nabla u(\bar{x}))$ for all $t\in[0,r]$. Suppose that $W(t)$ is a Jacobi field along $\bar{\gamma}$ with the property that $D_{t}W(0)=\sum_{j=1}^{n}(D^{2}u)(W(0),e_{j})\,e_{j}$, where $\{e_{1},\ldots,e_{n}\}$ denotes an orthonormal basis of $T_{\bar{x}}M$. If $W(\tau)=0$ for some $\tau\in(0,r)$, then $W$ vanishes identically.
Proof. We argue by contradiction. Suppose that $W(\tau)=0$ for some $\tau\in(0,r)$, but $W$ does not vanish identically. Using standard uniqueness results for ODE, we conclude that $D_{t}W(\tau)\neq 0$. Since $W$ is a Jacobi field, we obtain
$\int_{0}^{\tau}\big(|D_{t}W(t)|^{2}-R(\bar{\gamma}^{\prime}(t),W(t),\bar{\gamma}^{\prime}(t),W(t))\big)\,dt$
$=\langle D_{t}W(\tau),W(\tau)\rangle-\langle D_{t}W(0),W(0)\rangle$
$=-(D^{2}u)(W(0),W(0)).$
We define a vector field $\tilde{W}$ along $\bar{\gamma}$ by
$\tilde{W}(t)=\begin{cases}W(t)&\text{\rm for $t\in[0,\tau]$}\\ 0&\text{\rm for $t\in[\tau,r]$.}\end{cases}$
Clearly, $\tilde{W}(r)=0$ and
$\int_{0}^{r}\big(|D_{t}\tilde{W}(t)|^{2}-R(\bar{\gamma}^{\prime}(t),\tilde{W}(t),\bar{\gamma}^{\prime}(t),\tilde{W}(t))\big)\,dt$
$=-(D^{2}u)(\tilde{W}(0),\tilde{W}(0)).$
Since $D_{t}W(\tau)\neq 0$, the covariant derivative of $\tilde{W}(t)$ is discontinuous at the point $t=\tau$. Consequently, we can find a smooth vector field $Z$ along $\bar{\gamma}$ such that $Z(0)=\tilde{W}(0)$, $Z(r)=\tilde{W}(r)$, and
$\int_{0}^{r}\big(|D_{t}Z(t)|^{2}-R(\bar{\gamma}^{\prime}(t),Z(t),\bar{\gamma}^{\prime}(t),Z(t))\big)\,dt$
$<-(D^{2}u)(Z(0),Z(0)).$
This contradicts Lemma 5.5.
Proposition 5.7.
Assume that $r>0$ and $x\in A_{r}$. Then the function
$t\mapsto(1+t\,f(x)^{\frac{1}{n-1}})^{-n}\,|\det D\Phi_{t}(x)|$
is monotone decreasing for $t\in(0,r)$.
Proof. Let us fix a real number $r>0$ and a point $\bar{x}\in A_{r}$. We define $\bar{\gamma}(t):=\exp_{\bar{x}}(t\,\nabla u(\bar{x}))$ for all $t\in[0,r]$. Let $\{e_{1},\ldots,e_{n}\}$ be an orthonormal basis of $T_{\bar{x}}M$. For each $1\leq i\leq n$, we denote by $E_{i}(t)$ the unique parallel vector field along $\bar{\gamma}$ satisfying $E_{i}(0)=e_{i}$. Moreover, for each $1\leq i\leq n$, we denote by $X_{i}(t)$ the unique Jacobi field along $\bar{\gamma}$ satisfying $X_{i}(0)=e_{i}$ and
$D_{t}X_{i}(0)=\sum_{j=1}^{n}(D^{2}u)(e_{i},e_{j})\,e_{j}.$
It follows from Lemma 5.6 that $X_{1}(t),\ldots,X_{n}(t)$ are linearly independent for each $t\in(0,r)$.
For each $t\in[0,r]$, we define an $n\times n$-matrix $P(t)$ by
$P_{ij}(t)=\langle X_{i}(t),E_{j}(t)\rangle$
for $1\leq i,j\leq n$. Moreover, for each $t\in[0,r]$, we define an $n\times n$-matrix $S(t)$ by
$S_{ij}(t)=R(\bar{\gamma}^{\prime}(t),E_{i}(t),\bar{\gamma}^{\prime}(t),E_{j}(t))$
for $1\leq i,j\leq n$. The matrix $S(t)$ is symmetric for each $t\in[0,r]$. Moreover,
$\text{\rm tr}(S(t))=\text{\rm Ric}(\bar{\gamma}^{\prime}(t),\bar{\gamma}^{\prime}(t))\geq 0$
for each $t\in[0,r]$. Since the vector fields $X_{1}(t),\ldots,X_{n}(t)$ are Jacobi fields, we obtain
$P_{ij}^{\prime\prime}(t)$ $=\langle D_{t}D_{t}X_{i}(t),E_{j}(t)\rangle$
$=-R(\bar{\gamma}^{\prime}(t),X_{i}(t),\bar{\gamma}^{\prime}(t),E_{j}(t))$
$=-\sum_{k=1}^{n}P_{ik}(t)\,S_{kj}(t)$
for all $1\leq i,j\leq n$ and all $t\in[0,r]$. In other words, the function $P(t)$ satisfies $P^{\prime\prime}(t)=-P(t)S(t)$ for all $t\in[0,r]$. Moreover,
$P_{ij}(0)=\langle X_{i}(0),E_{j}(0)\rangle=\delta_{ij}$
and
$P_{ij}^{\prime}(0)=\langle D_{t}X_{i}(0),E_{j}(0)\rangle=(D^{2}u)(e_{i},e_{j})$
for $1\leq i,j\leq n$. In particular, the matrix $P^{\prime}(0)P(0)^{T}$ is symmetric. Moreover, the matrix
$\frac{d}{dt}(P^{\prime}(t)P(t)^{T})=P^{\prime\prime}(t)P(t)^{T}+P^{\prime}(t)P^{\prime}(t)^{T}=-P(t)S(t)P(t)^{T}+P^{\prime}(t)P^{\prime}(t)^{T}$
is symmetric for each $t\in[0,r]$. Consequently, the matrix $P^{\prime}(t)P(t)^{T}$ is symmetric for each $t\in[0,r]$.
Since $X_{1}(t),\ldots,X_{n}(t)$ are linearly independent for each $t\in(0,r)$, the matrix $P(t)$ is invertible for each $t\in(0,r)$. Since the matrix $P^{\prime}(t)P(t)^{T}$ is symmetric for each $t\in(0,r)$, it follows that the matrix $Q(t):=P(t)^{-1}P^{\prime}(t)$ is symmetric for each $t\in(0,r)$. The function $Q(t)$ satisfies the Riccati equation
$Q^{\prime}(t)=P(t)^{-1}P^{\prime\prime}(t)-P(t)^{-1}P^{\prime}(t)P(t)^{-1}P^{\prime}(t)=-S(t)-Q(t)^{2}$
for all $t\in(0,r)$. Moreover, since $Q(t)$ is symmetric, we obtain $\text{\rm tr}(Q(t)^{2})\geq\frac{1}{n}\,\text{\rm tr}(Q(t))^{2}$ for all $t\in(0,r)$. Since $\text{\rm tr}(S(t))\geq 0$, it follows that
$\frac{d}{dt}\text{\rm tr}(Q(t))=-\text{\rm tr}(S(t))-\text{\rm tr}(Q(t)^{2})\leq-\frac{1}{n}\,\text{\rm tr}(Q(t))^{2}$
for all $t\in(0,r)$. Using Lemma 5.3, we obtain
$\lim_{t\to 0}\text{\rm tr}(Q(t))=\Delta u(\bar{x})\leq n\,f(\bar{x})^{\frac{1}{n-1}}.$
A standard ODE comparison principle gives
$\text{\rm tr}(Q(t))\leq n\,\big(1+t\,f(\bar{x})^{\frac{1}{n-1}}\big)^{-1}\,f(\bar{x})^{\frac{1}{n-1}}$
for all $t\in(0,r)$.
We next observe that $\det P(t)>0$ if $t$ is sufficiently small. Since $P(t)$ is invertible for each $t\in(0,r)$, it follows that $\det P(t)>0$ for all $t\in(0,r)$. Moreover,
$\frac{d}{dt}\log\det P(t)=\text{\rm tr}(Q(t))\leq n\,\big(1+t\,f(\bar{x})^{\frac{1}{n-1}}\big)^{-1}\,f(\bar{x})^{\frac{1}{n-1}}$
for all $t\in(0,r)$. Consequently, the function
$t\mapsto(1+t\,f(\bar{x})^{\frac{1}{n-1}})^{-n}\,\det P(t)$
is monotone decreasing for $t\in(0,r)$.
Finally, the differential of $\Phi_{t}$ at the point $\bar{x}$ maps the vector $e_{i}\in T_{\bar{x}}M$ to the vector $X_{i}(t)\in T_{\bar{\gamma}(t)}M$. Consequently, $|\det D\Phi_{t}(\bar{x})|=\det P(t)$ for all $t\in(0,r)$. Putting these facts together, the assertion follows.
Corollary 5.8.
The Jacobian determinant of $\Phi_{r}$ satisfies
$|\det D\Phi_{r}(x)|\leq(1+r\,f(x)^{\frac{1}{n-1}})^{n}$
for each $r>0$ and each point $x\in A_{r}$.
Proof. Since $\lim_{t\to 0}|\det D\Phi_{t}(x)|=1$, the assertion follows from Proposition 5.7.
Using Lemma 5.4 and Corollary 5.8, we obtain
$\Big|\Big\{p\in M:\sup_{x\in D}d(x,p)<r\Big\}\Big|$ $\leq\int_{A_{r}}|\det D\Phi_{r}(x)|\,d\text{\rm vol}(x)$
$\leq\int_{D}(1+r\,f(x)^{\frac{1}{n-1}})^{n}\,d\text{\rm vol}(x)$
for all $r>0$. Finally, we divide by $r^{n}$ and send $r\to\infty$. This gives
$|B^{n}|\,\theta\leq\int_{D}f^{\frac{n}{n-1}}.$
Using the normalization condition (10), we conclude that
$\int_{D}|\nabla f|+\int_{\partial D}f=n\int_{D}f^{\frac{n}{n-1}}\geq n\,|B^{n}|^{\frac{1}{n}}\,\theta^{\frac{1}{n}}\,\Big(\int_{D}f^{\frac{n}{n-1}}\Big)^{\frac{n-1}{n}}.$
This completes the proof of Theorem 5.2.
6. The Fenchel-Willmore-Chen inequality for hypersurfaces in manifolds with nonnegative Ricci curvature
In this final section, we discuss a version of the Fenchel-Willmore-Chen inequality for hypersurfaces in manifolds with nonnegative Ricci curvature.
Theorem 6.1 (cf. [1], [22]).
Let $(M,g)$ be a complete noncompact manifold of dimension $n$ with nonnegative Ricci curvature. Let $\Sigma$ be a compact hypersurface in $M$ without boundary. Then
$\int_{\Sigma}\Big(\frac{|H|}{n-1}\Big)^{n-1}\geq|S^{n-1}|\,\theta,$
where $H$ denotes the mean curvature vector of $\Sigma$ and $\theta$ denotes the asymptotic volume ratio of $(M,g)$.
Theorem 6.1 is a consequence of an important estimate due to Heintze and Karcher [22]. Specifically, Theorem 2.1 in [22] gives an upper bound for the volume of a tubular neighborhood of $\Sigma$ of radius $r$. Theorem 6.1 follows from this inequality by sending $r\to\infty$; see the comment at the top of p. 452 in [22]. More recently, Agostiniani, Fogagnolo, and Mazzieri [1] gave an alternative proof of Theorem 6.1.
In the remainder of this section, we discuss the proof of Theorem 6.1, following the work of Heintze and Karcher.
We define
$U:=\{(x,y):x\in\Sigma,\,y\in T_{x}^{\perp}\Sigma,\,|y|<1\}.$
For each $r>0$, we denote by $A_{r}$ the set of all points $(\bar{x},\bar{y})\in U$ with the property that
$d\big(x,\exp_{\bar{x}}(r\bar{y})\big)\geq r\,|\bar{y}|$
for all $x\in\Sigma$. Moreover, for each $r>0$, we define a map $\Phi_{r}:U\to M$ by
$\Phi_{r}(x,y)=\exp_{x}(ry)$
for all $(x,y)\in U$.
Lemma 6.2.
For each $r>0$, the set
$\Big\{p\in M:\inf_{x\in\Sigma}d(x,p)<r\Big\}$
is contained in the image $\Phi_{r}(A_{r})$.
Proof. Let us fix a real number $r>0$. Let $p$ be a point in $M$ with the property that $\inf_{x\in\Sigma}d(x,p)<r$. We can find a point $\bar{x}\in\Sigma$ such that
$d(\bar{x},p)=\inf_{x\in\Sigma}d(x,p).$
Then $d(\bar{x},p)<r$. Let $\bar{\gamma}:[0,r]\to M$ be a minimizing geodesic such that $\bar{\gamma}(0)=\bar{x}$ and $\bar{\gamma}(r)=p$. The path $\bar{\gamma}$ minimizes the functional $\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=p$. The first variation formula implies that $\bar{\gamma}^{\prime}(0)\in T_{\bar{x}}^{\perp}\Sigma$. We define $\bar{y}=\bar{\gamma}^{\prime}(0)$. Then
$\exp_{\bar{x}}(r\bar{y})=\exp_{\bar{\gamma}(0)}(r\,\bar{\gamma}^{\prime}(0))=\bar{\gamma}(r)=p.$
Moreover,
$r\,|\bar{y}|=r\,|\bar{\gamma}^{\prime}(0)|=d(\bar{x},p).$
Since $d(\bar{x},p)<r$, it follows that $|\bar{y}|<1$. Therefore, $(\bar{x},\bar{y})\in U$. Finally, for each point $x\in\Sigma$, we have
$d\big(x,\exp_{\bar{x}}(r\bar{y})\big)=d(x,p)\geq d(\bar{x},p)=r\,|\bar{y}|.$
Thus, $(\bar{x},\bar{y})\in A_{r}$ and $\Phi_{r}(\bar{x},\bar{y})=p$. This completes the proof of Lemma 6.2.
Lemma 6.3.
Assume that $r>0$ and $(\bar{x},\bar{y})\in A_{r}$. We define $\bar{\gamma}(t):=\exp_{\bar{x}}(t\bar{y})$ for all $t\in[0,r]$. If $Z$ is a smooth vector field along $\bar{\gamma}$ such that $Z(0)\in T_{\bar{x}}\Sigma$ and $Z(r)=0$, then
$\langle I\!I(Z(0),Z(0)),\bar{y}\rangle+\int_{0}^{r}\big(|D_{t}Z(t)|^{2}-R(\bar{\gamma}^{\prime}(t),Z(t),\bar{\gamma}^{\prime}(t),Z(t))\big)\,dt\geq 0.$
Proof. Since $(\bar{x},\bar{y})\in A_{r}$, the path $\bar{\gamma}$ minimizes the functional $\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=\bar{\gamma}(r)$. Hence, the assertion follows from the second variation formula.
Lemma 6.4.
Assume that $r>0$ and $(\bar{x},\bar{y})\in A_{r}$. We define $\bar{\gamma}(t):=\exp_{\bar{x}}(t\bar{y})$ for all $t\in[0,r]$. Suppose that $W(t)$ is a Jacobi field along $\bar{\gamma}$ with the property that $W(0)\in T_{\bar{x}}\Sigma$ and $D_{t}W(0)=-\sum_{j=1}^{n-1}\langle I\!I(W(0),e_{j}),\bar{y}\rangle\,e_{j}$, where $\{e_{1},\ldots,e_{n-1}\}$ denotes an orthonormal basis of $T_{\bar{x}}\Sigma$. If $W(\tau)=0$ for some $\tau\in(0,r)$, then $W$ vanishes identically.
Proof. We argue by contradiction. Suppose that $W(\tau)=0$ for some $\tau\in(0,r)$, but $W$ does not vanish identically. Using standard uniqueness results for ODE, we conclude that $D_{t}W(\tau)\neq 0$. Since $W$ is a Jacobi field, we obtain
$\int_{0}^{\tau}\big(|D_{t}W(t)|^{2}-R(\bar{\gamma}^{\prime}(t),W(t),\bar{\gamma}^{\prime}(t),W(t))\big)\,dt$
$=\langle D_{t}W(\tau),W(\tau)\rangle-\langle D_{t}W(0),W(0)\rangle$
$=\langle I\!I(W(0),W(0)),\bar{y}\rangle.$
We define a vector field $\tilde{W}$ along $\bar{\gamma}$ by
$\tilde{W}(t)=\begin{cases}W(t)&\text{\rm for $t\in[0,\tau]$}\\ 0&\text{\rm for $t\in[\tau,r]$.}\end{cases}$
Clearly, $\tilde{W}(r)=0$ and
$\int_{0}^{r}\big(|D_{t}\tilde{W}(t)|^{2}-R(\bar{\gamma}^{\prime}(t),\tilde{W}(t),\bar{\gamma}^{\prime}(t),\tilde{W}(t))\big)\,dt$
$=\langle I\!I(\tilde{W}(0),\tilde{W}(0)),\bar{y}\rangle.$
Since $D_{t}W(\tau)\neq 0$, the covariant derivative of $\tilde{W}(t)$ is discontinuous at the point $t=\tau$. Consequently, we can find a smooth vector field $Z$ along $\bar{\gamma}$ such that $Z(0)=\tilde{W}(0)$, $Z(r)=\tilde{W}(r)$, and
$\int_{0}^{r}\big(|D_{t}Z(t)|^{2}-R(\bar{\gamma}^{\prime}(t),Z(t),\bar{\gamma}^{\prime}(t),Z(t))\big)\,dt$
$<\langle I\!I(Z(0),Z(0)),\bar{y}\rangle.$
This contradicts Lemma 5.5.
Proposition 6.5.
Assume that $r>0$ and $(x,y)\in A_{r}$. Then
$1-r\,\frac{\langle H(x),y\rangle}{n-1}\geq 0.$
Moreover, the function
$t\mapsto t^{-1}\,\Big(1-t\,\frac{\langle H(x),y\rangle}{n-1}\Big)^{-(n-1)}\,|\det D\Phi_{t}(x,y)|$
is monotone decreasing for $t\in(0,r)$.
Proof. Let us fix a real number $r>0$ and a point $(\bar{x},\bar{y})\in A_{r}$. We define $\bar{\gamma}(t):=\exp_{\bar{x}}(t\bar{y})$ for all $t\in[0,r]$. Let $\{e_{1},\ldots,e_{n-1}\}$ be an orthonormal basis of $T_{\bar{x}}\Sigma$. For each $1\leq i\leq n-1$, we denote by $E_{i}(t)$ the unique parallel vector field along $\bar{\gamma}$ satisfying $E_{i}(0)=e_{i}$. Moreover, for each $1\leq i\leq n-1$, we denote by $X_{i}(t)$ the unique Jacobi field along $\bar{\gamma}$ satisfying $X_{i}(0)=e_{i}$ and
$D_{t}X_{i}(0)=-\sum_{j=1}^{n-1}\langle I\!I(e_{i},e_{j}),\bar{y}\rangle\,e_{j}.$
For each $1\leq i\leq n-1$, $X_{i}(t)$ is a normal Jacobi field along $\bar{\gamma}$. In other words, $X_{i}(t)\in\text{\rm span}\{E_{1}(t),\ldots,E_{n-1}(t)\}$ for all $1\leq i\leq n-1$ and all $t\in[0,r]$. It follows from Lemma 6.4 that $X_{1}(t),\ldots,X_{n-1}(t)$ are linearly independent for each $t\in(0,r)$.
For each $t\in[0,r]$, we define an $(n-1)\times(n-1)$-matrix $P(t)$ by
$P_{ij}(t)=\langle X_{i}(t),E_{j}(t)\rangle$
for $1\leq i,j\leq n-1$. Moreover, for each $t\in[0,r]$, we define an $(n-1)\times(n-1)$-matrix $S(t)$ by
$S_{ij}(t)=R(\bar{\gamma}^{\prime}(t),E_{i}(t),\bar{\gamma}^{\prime}(t),E_{j}(t))$
for $1\leq i,j\leq n-1$. The matrix $S(t)$ is symmetric for each $t\in[0,r]$. Moreover,
$\text{\rm tr}(S(t))=\text{\rm Ric}(\bar{\gamma}^{\prime}(t),\bar{\gamma}^{\prime}(t))\geq 0$
for each $t\in[0,r]$. Since the vector fields $X_{1}(t),\ldots,X_{n-1}(t)$ are Jacobi fields, we obtain
$P_{ij}^{\prime\prime}(t)$ $=\langle D_{t}D_{t}X_{i}(t),E_{j}(t)\rangle$
$=-R(\bar{\gamma}^{\prime}(t),X_{i}(t),\bar{\gamma}^{\prime}(t),E_{j}(t))$
$=-\sum_{k=1}^{n-1}P_{ik}(t)\,S_{kj}(t)$
for all $1\leq i,j\leq n-1$ and all $t\in[0,r]$. In other words, the function $P(t)$ satisfies $P^{\prime\prime}(t)=-P(t)S(t)$ for all $t\in[0,r]$. Moreover,
$P_{ij}(0)=\langle X_{i}(0),E_{j}(0)\rangle=\delta_{ij}$
and
$P_{ij}^{\prime}(0)=\langle D_{t}X_{i}(0),E_{j}(0)\rangle=-\langle I\!I(e_{i},e_{j}),\bar{y}\rangle$
for $1\leq i,j\leq n-1$. In particular, the matrix $P^{\prime}(0)P(0)^{T}$ is symmetric. Moreover, the matrix
$\frac{d}{dt}(P^{\prime}(t)P(t)^{T})=P^{\prime\prime}(t)P(t)^{T}+P^{\prime}(t)P^{\prime}(t)^{T}=-P(t)S(t)P(t)^{T}+P^{\prime}(t)P^{\prime}(t)^{T}$
is symmetric for each $t\in[0,r]$. Consequently, the matrix $P^{\prime}(t)P(t)^{T}$ is symmetric for each $t\in[0,r]$.
Since $X_{1}(t),\ldots,X_{n-1}(t)$ are linearly independent for each $t\in(0,r)$, the matrix $P(t)$ is invertible for each $t\in(0,r)$. Since the matrix $P^{\prime}(t)P(t)^{T}$ is symmetric for each $t\in(0,r)$, it follows that the matrix $Q(t):=P(t)^{-1}P^{\prime}(t)$ is symmetric for each $t\in(0,r)$. The function $Q(t)$ satisfies the Riccati equation
$Q^{\prime}(t)=P(t)^{-1}P^{\prime\prime}(t)-P(t)^{-1}P^{\prime}(t)P(t)^{-1}P^{\prime}(t)=-S(t)-Q(t)^{2}$
for all $t\in(0,r)$. Moreover, since $Q(t)$ is symmetric, we obtain $\text{\rm tr}(Q(t)^{2})\geq\frac{1}{n-1}\,\text{\rm tr}(Q(t))^{2}$ for all $t\in(0,r)$. Since $\text{\rm tr}(S(t))\geq 0$, it follows that
$\frac{d}{dt}\text{\rm tr}(Q(t))=-\text{\rm tr}(S(t))-\text{\rm tr}(Q(t)^{2})\leq-\frac{1}{n-1}\,\text{\rm tr}(Q(t))^{2}$
for all $t\in(0,r)$. Moreover,
$\lim_{t\to 0}\text{\rm tr}(Q(t))=-\langle H(\bar{x}),\bar{y}\rangle.$
A standard ODE comparison principle implies that
$1-r\,\frac{\langle H(\bar{x}),\bar{y}\rangle}{n-1}\geq 0$
and
$\text{\rm tr}(Q(t))\leq-\Big(1-t\,\frac{\langle H(\bar{x}),\bar{y}\rangle}{n-1}\Big)^{-1}\,\langle H(\bar{x}),\bar{y}\rangle$
for all $t\in(0,r)$.
We next observe that $\det P(t)>0$ if $t$ is sufficiently small. Since $P(t)$ is invertible for each $t\in(0,r)$, it follows that $\det P(t)>0$ for all $t\in(0,r)$. Moreover,
$\frac{d}{dt}\log\det P(t)=\text{\rm tr}(Q(t))\leq-\Big(1-t\,\frac{\langle H(\bar{x}),\bar{y}\rangle}{n-1}\Big)^{-1}\,\langle H(\bar{x}),\bar{y}\rangle$
for all $t\in(0,r)$. Consequently, the function
$t\mapsto\Big(1-t\,\frac{\langle H(\bar{x}),\bar{y}\rangle}{n-1}\Big)^{-(n-1)}\,\det P(t)$
is monotone decreasing for $t\in(0,r)$.
Finally, it is easy to see that $|\det D\Phi_{t}(\bar{x},\bar{y})|=t\,\det P(t)$ for all $t\in(0,r)$. Putting these facts together, the assertion follows.
Corollary 6.6.
The Jacobian determinant of $\Phi_{r}$ satisfies
$|\det D\Phi_{r}(x,y)|\leq r\,\Big(1-r\,\frac{\langle H(x),y\rangle}{n-1}\Big)^{n-1}$
for each $r>0$ and each point $(x,y)\in A_{r}$.
Proof. Since $\lim_{t\to 0}t^{-1}\,|\det D\Phi_{t}(x,y)|=1$, the assertion follows from Proposition 6.5.
Using Lemma 6.2 and Corollary 6.6, we obtain
$\Big|\Big\{p\in M:\inf_{x\in\Sigma}d(x,p)<r\Big\}\Big|$
$\leq\int_{\Sigma}\bigg(\int_{\{y\in T_{x}^{\perp}\Sigma:|y|<1\}}|\det D\Phi_{r}(x,y)|\,1_{A_{r}}(x,y)\,dy\bigg)\,d\text{\rm vol}(x)$
$\leq\int_{\Sigma}\bigg(\int_{\{y\in T_{x}^{\perp}\Sigma:|y|<1\}}r\,\Big(1-r\,\frac{\langle H(x),y\rangle}{n-1}\Big)_{+}^{n-1}\,dy\bigg)\,d\text{\rm vol}(x)$
for all $r>0$. We divide by $r^{n}$ and send $r\to\infty$. This gives
$|B^{n}|\,\theta\leq\int_{\Sigma}\bigg(\int_{\{y\in T_{x}^{\perp}\Sigma:|y|<1\}}\Big(-\frac{\langle H(x),y\rangle}{n-1}\Big)_{+}^{n-1}\,dy\bigg)\,d\text{\rm vol}(x).$
On the other hand,
$\int_{\{y\in T_{x}^{\perp}\Sigma:|y|<1\}}(-\langle H(x),y\rangle)_{+}^{n-1}\,dy=\frac{1}{n}\,|H(x)|^{n-1}=\frac{|B^{n}|}{|S^{n-1}|}\,|H(x)|^{n-1}$
for each point $x\in\Sigma$. Putting these facts together, we conclude that
$|B^{n}|\,\theta\leq\frac{|B^{n}|}{|S^{n-1}|}\int_{\Sigma}\Big(\frac{|H(x)|}{n-1}\Big)^{n-1}\,d\text{\rm vol}(x).$
This completes the proof of Theorem 5.2.